I need help completing the two chart. Please and thank you if you show your work.

Data Table 1: Preparation and Standardization of a Sodium Hydroxide Solution Parameter Trial #1 | Trial #2 Trial #3 Wt. of KHP bottle before sample taken (w) Wt. of KHP bottle after sample taken (w2) Wt. of KHP (w = w1-w2) Initial buret reading (V Final buret reading (V2) Volume of NaOH delivered (V = V2-V1) Number of total drops delivered Number of drops delivered at equivalence point | Volume of NaOH delivered at equivalence point | Molarity of NaOH Average Molarity of NaOH (Mave Absolute Deviation (M Mave Average Deviation (AD) Relative Average Deviation (RAD), ppt 0, 230 0.3 n C 2. ンcha, Iyo Yala 73 Relative Average Deviaton (RAD) ppt RADMe × 1000
Answer #1

For data table 1 -

Absolute deviation -

For trial 1 = 0.06306 - 0.07435

= - 0.01129

For trial 2 = 0.07522 - 0.07435

= 0.00087

For trial 3 = 0.08476 - 0.07435

= 0.01041

Average deviation -

AD = - 0.01129 + 0.00087 + 0.01041 / 3

= - 0.00001

Relative average deviation-

RAD = AD / Mave * 1000

= - 0.00001 / 0.07435 * 1000

= -0.13449